Please help..i am at a loss here.
Solve...
8x + 9 = 25
4x + 10 = 30
Solve using the principles together
3x - 18 = 9 The solution is x = ?

Answer by HyperBrain(694)

8x+9=25
8x+9-9=25-9
8x+0=16
8x=16
x=2

4x+10=30
4x+10-10=30-10
4x+0=20
4x=20
x=5
3x-18=9
3x-18+18=9+18
3x+0=27
3x=27
x=9
Power up,
HyperBrain!